# Music-Aligned Holistic 3D Dance Generation via Hierarchical Motion Modeling

### v2

@@ Abstract
Figure 1: We introduce SoulDance, a high-quality, comprehensive dance dataset that incorporates body motions, hand gestures, and facial expressions. The dataset processing pipeline (left) consists of separate steps for capturing facial expressions and body-hand movements. Moreover, we present SoulNet, the first framework able to generate expressive, holistic dance, as demonstrated in the results (right).
**footnotetext: This work is done during the internship at ByteDance\$$\dagger$\$$\dagger$\$footnotetext: Corresponding author
@@ 2 Related Work
Dataset Pos/Rot Joints num Hand Face Genres Mocap Single Dance Group Dance SMPL Total hours avg Sec per Seq
@@
Table 1: Comparisons of 3D Dance Datasets. “Pos" and “Rot" represent 3D position and rotation information, respectively. “52" refers to the parameters associated with ARKit blendshapes. “avg Sec per Seq" indicates the average duration in seconds for each sequence.
@@ 4.1 Hierarchical Residual Vector Quantization
where $sg[\cdot]$ denotes the stop-gradient operation. The parameters $\alpha$, $\beta$, and $\gamma$ are factors for the embedding constraints.
@@ 4.2 Music-Aligned Generative Model
$\lambda_{b}$ and $\lambda_{w}$ are hyperparameters, both set to 0.5. The dance-music alignment priors for both $\mathcal{L}_{\text{Align-whole}}$ and $\mathcal{L}_{\text{Align-body}}$ are derived from the Music-Motion Retrieval Module.
@@ 4.3 Music-Motion Retrieval Module
To address this limitation, we propose the Music-Motion Retrieval (MMR) module, a cross-modal alignment framework designed specifically for the music-dance domain. As illustrated in Figure 3 (right), our architecture consists of three principal components: (1) Motion Encoding, where input motion sequences $M\in\mathbb{R}^{T\times D_{m}}$ are compressed into latent codes $\mathbf{z}=\{z_{i}\}_{i=1}^{d}$ through a temporal encoder $\mathcal{E}_{\text{motion}}$ with learnable downsampling operations. (2) Music Processing, in which raw audio inputs $C$ are first encoded using the pretrained Jukebox encoder [8], followed by a dedicated music encoder $\mathcal{E}_{\text{music}}$ that projects features into a motion-aligned latent space $\mathbf{c}=\{m_{i}\}_{i=1}^{d}$. (3) Cross-Modal Fusion, where the aligned latent representations $\mathbf{z}$ and $\mathbf{c}$ are concatenated and decoded through a transformer-based motion decoder $\mathcal{D}_{\text{motion}}$ to reconstruct coherent dance sequences. Our encoder-decoder architecture extends the TEMOS framework [42] with modality-specific adaptations. The MMR module is optimized through two complementary objectives: a Reconstruction Loss and a Contrastive Learning objective using InfoNCE loss [39] to maximize mutual information between positive music-motion pairs while minimizing similarity for negative pairs. We apply $\mathcal{L}_{\text{Align-body}}$ to guide the generation of bodily structure, and employ $\mathcal{L}_{\text{Align-whole}}$ to align the whole-body motion with musical affect. See Appendix H for more details.
During inference, the music signal $C$ and motion sequences $M\in\mathbb{R}^{T\times D_{m}}$ are encoded into latent representations $\mathbf{z}$ and $\mathbf{c}$ via the MMR encoder. These representations are then utilized to compute alignment losses $\mathcal{L}_{\text{Align-body}}$ and $\mathcal{L}_{\text{Align-whole}}$ through the InfoNCE objective [39]. In the MAGM dance generation framework, $\mathcal{L}_{\text{Align-body}}$ enforces local temporal alignment between body movements and musical beats, while $\mathcal{L}_{\text{Align-whole}}$ regulates global synchronization of holistic motion dynamics with musical affect. To achieve this dual granularity, we train two specialized MMR modules using two different datasets. Further details are provided in Appendix H.
@@ 5.1 Comparison to SOTA Methods
Public benchmarks. (1) FID score. Fréchet inception distance (FID) is widely used to measure how close the distribution of the generated dances is to that of the ground truth. We follow the approach introduced in HumanML3D [17], which is widely adopted in motion generation studies for assessing motion quality [19, 38, 18, 32]. (2) Diversity. Diversity evaluate the average feature distance between generated dances for different input music. The same feature extractor used in FID is used again. (3) Hand FID score and Hand Diversity. Similarly, we extract hand motion features and calculate the FID and diversity for hand motion. (4) Multimodality. We follow Guo et al.[17] to evaluate the average feature distance between the 10 choreography versions of every music. This metric measures the model’s ability to generate different dances for the same music. (5) Run Time. We evaluated the average runtime required for the model to generate dance sequences of equal length during the inference process. Qualitative results further demonstrate that our method, SoulNet, achieves high-quality and diverse holistic dance generation. Please refer to Appendix F.
@@ 5.2 Ablation Study
Effect of Quantization Layers. In Table 3, we analyze the impact of different quantization layer counts $V$ for both HRVQ and RVQ methods. Generally, adding more residual VQ layers enhances reconstruction precision, though this also increases the computational load on the MAGM for dance token generation. As shown in Figure 4, we observe a noticeable decline in generation quality when the number of residual layers exceeds 5 or 6, both in RVQ and HRVQ. Therefore, to balance reconstruction accuracy and generation efficiency, SoulNet is designed with $V=5$ (i.e., $5$ residual VQ layers) for HRVQ.
@@ 6 Conclusion
In this paper, we introduce SoulDance, a large-scale, high-quality 3D holistic dance dataset for music-driven dance generation, capturing synchronized body movements, hand gestures, and facial expressions. We further present SoulNet, which utilizes HRVQ for joint modeling and efficient quantization of holistic dance. Using the music-dance alignment prior from MMR to supervise MAGM, SoulNet generates expressive holistic dance sequences that are aligned with the input music. Additionally, we propose two novel evaluation metrics: EmotionAlign and the MMR-Matching Score, to assess the alignment of facial and body motions with music. Both quantitative and qualitative results demonstrate that SoulNet achieves SOTA performance in generating holistic dance aligned with music. However, our approach does not consider the generation of elements beyond the human body, such as the coordination of clothing and props in the dance performance. Addressing these limitations will be a key direction for future research.
@@ Acknowledgements
We sincerely thank the anonymous ICCV reviewers for their invaluable feedback and suggestions. We would like to thank ByteDance for sponsoring the research. Xiaojie Li would also like to thank his colleagues in the lab, Weixiang Zhang, Rongwei Lu, Jiajun Luo and Duo Wu, for their constructive comments on improving the quality of this paper. This work was supported in part by National Key Research and Development Project of China (Grant No. 2023YFF0905502), National Natural Science Foundation of China (Grant No. 92467204, 62472249), National Key R&D Program of China (No. 2022YFF0902303), and Shenzhen Science and Technology Program (Grant No. JCYJ20220818101014030 and KJZD20240903102300001).
@@ Appendix E Dance Reconstruction Evaluation Metrics
where $v_{j}$ represents the ground truth positions of the facial vertices, and $\tilde{v}_{j}$ denotes the corresponding vertices reconstructed by HRVQ. The metric is averaged over $N$ frames to provide a comprehensive evaluation of facial reconstruction quality.
@@ Appendix F Additional Qualitative Results
Comparison with SOTA Methods. SoulNet demonstrates exceptional qualitative performance on both the AIST++ and SoulDance datasets. In Figure 10, FACT [31] generates dance sequences where, after the initial two seconds, body and hand movements become mostly static, and facial expressions are entirely absent. EDGE [53] produces convincing body movements but often fails to generate detailed hand motions and lacks expressive facial output. Bailando [50] captures body movements and facial expressions effectively but suffers from joint dislocations during turns and inadequate hand generation. FineNet [32] delivers satisfactory dance and hand motions but struggles with fine hand articulation and facial expressiveness. In contrast, SoulNet not only generates diverse and dynamic body movements but also excels in capturing intricate hand and facial details. As shown in Figure 9, on the AIST++ dataset, SoulNet achieves superior alignment with the musical beat and demonstrates greater diversity in generated dance sequences compared to other methods. Videos are available on our website.
@@ Appendix H Training: Music-Motion Retrieval Module
Dataset. To establish robust dance-music alignment priors, we gathered high-quality open-source music-dance datasets: AIST++ [31], Finedance [32], PhantomDance [30], and SoulDance (totaling 25 hours). Following the preprocessing protocol of EDGE [53] with default temporal segmentation parameters, we derive two specialized motion representations for training distinct Music-Motion Retrieval modules. For Body-Alignment MMR, all sequences are processed using the HumanML3D [16] motion representation ($D_{m}=263$). For Whole-Alignment MMR, SoulDance dataset is reformatted via our Holistic Dance Representation (Appendix D, $D_{m}=723$), encoding holistic motion movements. Finally, all datasets are partitioned into training/validation/test splits (8:1:1 ratio) using a stratified strategy that preserves music genre and dance style distributions. Crucially, we enforce temporal alignment constraints between SoulNet and MMR training subsets within AIST++ and SoulDance to prevent data leakage—ensuring no overlapping music clips or motion segments exist across models. The key distinctions between $\mathcal{L}_{\text{Align-body}}$ and $\mathcal{L}_{\text{Align-whole}}$ are: (a) $\mathcal{L}_{\text{Align-body}}$ enforces local temporal alignment between body movement and musical rhythm, whereas $\mathcal{L}_{\text{Align-whole}}$ captures global motion-music synchronization including emotional cues. (b) Two specialized MMR modules, $\text{MMR}_{\text{body}}$ and $\text{MMR}_{\text{whole}}$, are pre-trained to provide supervisory signals for the respective losses. $\text{MMR}_{\text{body}}$ is trained on body-only motion data from public datasets (AIST++ [31], FineDance [32]], PhantomDance [30]) using a 263-dimensional motion representation ($D_{m}=263$). In contrast, $\text{MMR}_{\text{whole}}$ is trained on a subset of SoulDance holistic motion data (including body, hands and face) with a 723-dimensional representation ($D_{m}=723$).
@@
Experiments. We conducted both qualitative and quantitative experiments to evaluate the performance of the MMR module. As shown in Table 8, MMR demonstrates exceptional retrieval capabilities. Visualized retrieval results further validate this observation. For comparison, we provide two music samples, each paired with two high-similarity dance sequences and two low-similarity dance sequences. Figure 12 and the demo examples (available on our project website) illustrate that our retrieval results align better with the beat and rhythm. In these examples, higher similarity scores indicate a stronger correlation between the music and the retrieved dance motions.

### v3

@@ Abstract
Figure 1: We introduce SoulDance, a high-quality, comprehensive dance dataset that incorporates body motions, hand gestures, and facial expressions. The dataset processing pipeline (left) consists of separate steps for capturing facial expressions and body-hand movements. Moreover, we present SoulNet, the first framework able to generate expressive and holistic dance, as demonstrated in the results (right).
**footnotetext: This work was partially conducted during an internship at ByteDance.$\dagger$\$$\dagger$\$footnotetext: Corresponding author.
@@ 2 Related Work
Dataset Pos/Rot Joints num Hand Face Genres Mocap Single Dance Group Dance SMPL
Total
hours
avg Sec
per Seq
@@
Table 1: Comparisons of 3D Dance Datasets. “Pos” and “Rot” represent 3D position and rotation information, respectively. “52” refers to the parameters associated with ARKit blendshapes. “avg Sec per Seq” indicates the average duration in seconds for each sequence.
@@ 4.1 Hierarchical Residual Vector Quantization
where $sg[\cdot]$ denotes the stop-gradient operation. The parameters $\alpha$, $\beta$, and $\gamma$ are factors for the embedding constraints. Since HRVQ relies on discrete VQ token sampling, we adopt the Gumbel-Softmax [21] trick to enable gradient backpropagation.
@@ 4.2 Music-Aligned Generative Model
$\lambda_{b}$ and $\lambda_{w}$ are hyperparameters, both set to 0.5. $\hat{T}_{0}$ alone is insufficient to generate fine-grained dance, so we use $V$ residual layers to predict the residuals $\hat{T}_{1:V}$. These are concatenated with $\hat{T}_{0}$ to form $\hat{T}_{0:V}$, which is decoded by $\mathcal{D}_{\text{whole}}$ to generate the final motion. The dance-music alignment priors for both $\mathcal{L}_{\text{Align-whole}}$ and $\mathcal{L}_{\text{Align-body}}$ are derived from the Music-Motion Retrieval Module.
@@ 4.3 Music-Motion Retrieval Module
To address this limitation, we propose the Music-Motion Retrieval (MMR) module, a cross-modal alignment framework designed specifically for the music-dance domain. As illustrated in Figure 3 (right), our architecture consists of three principal components: (1) Motion Encoding, where input motion sequences $M\in\mathbb{R}^{T\times D_{m}}$ are compressed into latent codes $\mathbf{z}=\{z_{i}\}_{i=1}^{d}$ through a temporal encoder $\mathcal{E}_{\text{motion}}$ with learnable downsampling operations. (2) Music Processing, in which raw audio inputs $C$ are first encoded using the pretrained Jukebox encoder [8], followed by a dedicated music encoder $\mathcal{E}_{\text{music}}$ that projects features into a motion-aligned latent space $\mathbf{c}=\{m_{i}\}_{i=1}^{d}$. (3) Cross-Modal Fusion, where the aligned latent representations $\mathbf{z}$ and $\mathbf{c}$ are concatenated and decoded through a transformer-based motion decoder $\mathcal{D}_{\text{motion}}$ to reconstruct coherent dance sequences. Our encoder-decoder architecture extends the TEMOS framework [41] with modality-specific adaptations. We apply $\mathcal{L}_{\text{Align-body}}$ to guide the generation of bodily structure, and employ $\mathcal{L}_{\text{Align-whole}}$ to align the whole-body motion with musical affect. Both losses are derived from an alignment objective:
$\small\begin{split}\mathcal{L}_{\text{Align}}&=-\frac{1}{2N}\sum_{i}\left(\log\frac{e^{S_{ii}}}{\sum_{j}e^{S_{ij}}}+\log\frac{e^{S_{ii}}}{\sum_{j}e^{S_{ji}}}\right)~,\end{split}$ (8)
The difference arises from the similarity matrix $S$, as it is derived from different motion representations.
During inference, the music signal $C$ and motion sequences $M\in\mathbb{R}^{T\times D_{m}}$ are encoded into latent representations $\mathbf{z}$ and $\mathbf{c}$ via the MMR encoder. In the MAGM dance generation framework, $\mathcal{L}_{\text{Align-body}}$ enforces local temporal alignment between body movements and musical beats, while $\mathcal{L}_{\text{Align-whole}}$ regulates global synchronization of holistic motion dynamics with musical affect. To achieve this dual granularity, we train two specialized MMR modules using two different datasets. Please refer to Appendix H.
@@ 5.1 Comparison to SOTA Methods
Public benchmarks. (1) FID score. Fréchet inception distance (FID) is widely used to measure how close the distribution of the generated dances is to that of the ground truth. We follow the approach introduced in HumanML3D [17], which is widely adopted in motion generation studies for assessing motion quality [19, 38, 18, 32]. (2) Diversity. Diversity evaluate the average feature distance between generated dances for different input music. The same feature extractor used in FID is used again. (3) Hand FID score and Hand Diversity. Similarly, we extract hand motion features and calculate the FID and diversity for hand motion. (4) Multimodality. We follow Guo et al.[17] to evaluate the average feature distance between the 10 choreography versions of every music. This metric measures the model’s ability to generate different dances for the same music. (5) Run Time. We evaluated the average runtime to generate dance sequences of equal length during the inference process. Qualitative results further demonstrate that our method, SoulNet, achieves high-quality and diverse holistic dance generation. Please refer to Appendix F.
@@ 5.2 Ablation Study
Effect of Quantization Layers. In Table 3, we analyze the impact of different quantization layer counts $V$ for both HRVQ and RVQ methods. Generally, adding more residual VQ layers enhances reconstruction precision, though this also increases the computational load on the MAGM for dance token generation. As shown in Figure 4, we observe a noticeable decline in generation quality when the number of residual layers exceeds 5 or 6, both in RVQ and HRVQ. Therefore, to balance reconstruction accuracy and generation efficiency, SoulNet is designed with $V=5$ (i.e., 555 residual VQ layers) for HRVQ.
@@ 6 Conclusion
In this paper, we introduce SoulDance, a large-scale, high-quality 3D holistic dance dataset for music-driven dance generation, capturing synchronized body movements, hand gestures, and facial expressions. We further present SoulNet, which utilizes HRVQ for joint modeling and efficient quantization of holistic dance. Using the music-dance alignment prior from MMR to supervise MAGM, SoulNet generates expressive holistic dance sequences that are aligned with the input music. Additionally, we propose two novel evaluation metrics: EmotionAlign and the MMR-Matching Score, to assess the alignment of facial and body motions with music. Both quantitative and qualitative results demonstrate that SoulNet achieves SOTA performance in generating holistic dance aligned with music.
@@ Acknowledgements
We sincerely thank the anonymous ICCV reviewers for their invaluable feedback and suggestions. Xiaojie Li would also like to thank his colleagues in the lab, Weixiang Zhang, Rongwei Lu, Jiajun Luo and Duo Wu, for their constructive comments on improving the quality of this paper. This work was supported in part by National Key Research and Development Project of China (Grant No. 2023YFF0905502), National Natural Science Foundation of China (Grant No. 92467204, 62472249), National Key R&D Program of China (No. 2022YFF0902303), and Shenzhen Science and Technology Program (Grant No. JCYJ20220818101014030 and KJZD20240903102300001).
@@ Appendix E Dance Reconstruction Evaluation Metrics
where $v_{j}$ represents the ground truth positions of the facial vertices, and $\tilde{v}_{j}$ denotes the corresponding vertices reconstructed by HRVQ. The metric is averaged over $N$ frames to evaluate facial reconstruction quality.
@@ Appendix F Additional Qualitative Results
Comparison with SOTA Methods. SoulNet demonstrates exceptional qualitative performance on both the AIST++ and SoulDance datasets. In Figure 10, FACT [31] generates dance sequences where, after the initial two seconds, body and hand movements become mostly static, and facial expressions are entirely absent. EDGE [52] produces convincing body movements but often fails to generate detailed hand motions and lacks expressive facial output. Bailando [49] captures body movements and facial expressions effectively but suffers from joint dislocations during turns and inadequate hand generation. FineNet [32] delivers satisfactory dance and hand motions but struggles with fine hand articulation and facial expressiveness. In contrast, SoulNet not only generates diverse and dynamic body movements but also excels in capturing intricate hand and facial details. As shown in Figure 9, on the AIST++ dataset, SoulNet achieves superior alignment with the musical beat and demonstrates greater diversity in generated dance sequences compared to other methods.
@@ Appendix H Training: Music-Motion Retrieval Module
Dataset. To establish robust dance-music alignment priors, we gathered high-quality open-source music-dance datasets: AIST++ [31], Finedance [32], PhantomDance [30], and SoulDance (totaling 25 hours). Following the preprocessing protocol of EDGE [52] with default temporal segmentation parameters, we derive two specialized motion representations for training distinct Music-Motion Retrieval modules. For Body-Alignment MMR, all sequences are processed using the HumanML3D [16] motion representation ($D_{m}=263$). For Whole-Alignment MMR, SoulDance dataset is reformatted via our Holistic Dance Representation (Appendix D, $D_{m}=723$), encoding holistic motion movements. Finally, all datasets are partitioned into training/validation/test splits (8:1:1 ratio) using a stratified strategy that preserves music genre and dance style distributions.
Technical Details. Crucially, we enforce temporal alignment constraints between SoulNet and MMR training subsets within AIST++ and SoulDance to prevent data leakage—ensuring no overlapping music clips or motion segments exist across models. Two specialized MMR modules, $\text{MMR}_{\text{body}}$ and $\text{MMR}_{\text{whole}}$, are pre-trained to provide supervisory signals for the respective losses. $\text{MMR}_{\text{body}}$ is trained on body-only motion data from public datasets (AIST++ [31], FineDance [32]], PhantomDance [30]) using a 263-dimensional motion representation ($D_{m}=263$). In contrast, $\text{MMR}_{\text{whole}}$ is trained on a subset of SoulDance holistic motion data (including body, hands and face) with a 723-dimensional representation ($D_{m}=723$).
@@
Experiments. We conducted both qualitative and quantitative experiments to evaluate the performance of the MMR module. As shown in Table 8, MMR demonstrates exceptional retrieval capabilities. Visualized retrieval results further validate this observation. For comparison, we provide two music samples, each paired with two high-similarity dance sequences and two low-similarity dance sequences. Figure 12 and the demo examples illustrate that our retrieval results align better with the beat and rhythm. In these examples, higher similarity scores indicate a stronger correlation between the music and the retrieved dance motions.